# IUT-Plug: A Plug-in tool for Interleaved Image-Text Generation

### v1

@@ 1 Introduction
Despite remarkable advancements, modern Vision Language Models (VLMs) suffer from a fundamental limitation known as multimodal context drift. Multimodal context drift refers to a progressive loss of cross-modal consistency during extended interleaved image-text interactions. Although VLMs excel at interpreting static image-text pairs (Rombach et al., 2022; Ramesh et al., 2022), they often fail to maintain coherence across multi-turn sequences. Specifically, (1) Logic drift occurs when the joint semantic content derived from the input image and accompanying text is not faithfully reflected in either the textual answer or the subsequently generated image, leading to contradictions between perception, instruction, and output. (2) Entity identity drift manifests when concrete referents—such as named characters or objects present in the input image or text—are misidentified, lose their attributes, or disappear entirely in later generations. (3) Style drift happens when distinctive visual characteristics of the input image, including artistic medium, color palette, lighting, or composition, are not preserved in the output image despite being explicitly conditioned on that input (Malakouti & Kovashka, 2025; Lorenz et al., 2024; Goyal et al., 2021). These three forms of drift reveal a systemic failure in current VLMs to maintain a unified, persistent representation of the multimodal scene across turns (Bougzime et al., 2025; Marcus, 2020).
@@
To bridge this gap, we propose IUT-Plug, a lightweight and model-agnostic plug-in module. IUT-Plug is built around Image Understanding Trees (IUTs), which are hierarchical symbolic structures. These structures explicitly represent the entities in a visual scene, their attributes, and the relationships among them. The module integrates seamlessly with existing large vision-language models without requiring costly end-to-end retraining (Li et al., 2024; Chen et al., 2024a). It operates within the neuro-symbolic paradigm (Kautz, 2022; Sarker et al., 2021; d’Avila Garcez & Lamb, 2019). By design, IUT-Plug separates the perceptual strength of neural networks from the logical precision of symbolic reasoning. It uses the IUT as a factual scaffold to guide both text and image generation (Ferreira et al., 2023; Pan et al., 2023). This combination allows the system to maintain consistency while preserving the flexibility of modern generative models.
@@
IUT-Plug bridges perception and generation through a shared, explicit world model. In the perception phase, the input image is parsed into an Image Understanding Tree (IUT), which represents entities, their attributes, and their relationships in a structured symbolic format. As new interleaved image-text instructions arrive, this state is incrementally updated to reflect evolving semantics. The refined IUT then actively guides both the language response and the prompt for the downstream text-to-image generator. By anchoring the entire generation process in this symbolic scaffold, IUT-Plug ensures that logical coherence, entity identity, and visual style are faithfully propagated from the input to all subsequent outputs. This design directly addresses the three core forms of multimodal context drift that undermine current interleaved VLM pipelines.
@@ 2.1 Interleaved Vision-Language Models
The ability to process and generate sequences of interleaved images and text marks a major advance over single turn multimodal systems. Pioneering architectures like Flamingo (Alayrac et al., 2022) introduced gated cross-attention to ingest multimodal streams. More recent systems have further enhanced generative fluency by integrating vision-language models (VLMs) with text-to-image (T2I) generators, such as MiniGPT-5 (Zheng et al., 2023), Emu2 (Sun et al., 2024), and MM-Interleaved (Chen et al., 2024b). They often using visual tokens or feature synchronizers to bridge modalities (Wu et al., 2023b). However, no explicit mechanism exists to propagate the original visual context—especially its logic, entity identity, and style. As a result, the system suffers from multimodal context drift. The joint semantics of the input are progressively lost or distorted across turns.
@@ 2.2 Image Generation from Structured Representations
Structured symbolic representations have long been used to guide image synthesis. Much of this work centers on scene graphs (Xu et al., 2022). Models such as SG2IM (Johnson et al., 2018) and SGDiff (Yang et al., 2022) show that a static graph—encoding objects, their attributes, and pairwise relationships—can be effectively translated into a coherent image. The inverse task is addressed by Scene Graph Generation (SGG) models (Krishna et al., 2017), which parse a single image into such a graph. However, these approaches treat the structured representation as a one-time input or output. They do not support updates across interactions. As a result, they are inherently limited to single-turn generation. Multi-turn consistency remains out of reach in this paradigm.
@@ 2.3 Neuro-Symbolic Generative Models
Neuro-symbolic AI aims to combine the perceptual strength of neural networks with the reasoning power of symbolic systems (Marcus, 2020; d’Avila Garcez & Lamb, 2019). In generative modeling, several methods have explored symbolic guidance. Neuro-Symbolic Diffusion (NSD) (Christopher et al., 2025) injects logical constraints directly into the denoising steps of diffusion models. Control-GPT (Wu et al., 2023a) uses programmatic sketches to control layout and composition. These approaches often rely on hard constraints or procedural specifications. While effective for constrained tasks, they can limit the generative flexibility of the underlying model. Moreover, they typically assume a fixed symbolic specification at the start of generation. They lack mechanisms to maintain or evolve a world state over extended, open-ended interactions.
@@ 3 Evaluation Framework
Figure 2: Overview of our evaluation metric. The evaluation model is fine-tuned on 3,000 sample data annotated by experts. For each question-answer pair, we use GPT-5 to generate three dynamic evaluation criteria, and then employ the evaluation model to output “yes” or “no” for each criterion.
@@
Seen as Figure 3. For each input-output tuple $(Q_{t},I_{t},A_{t})$, where $Q_{t}\in\mathcal{Q}$ is a natural language instruction, $I_{t}\in\mathbb{R}^{H\times W\times 3}$ is the reference image, and $A_{t}=(T_{t},I_{t}^{\prime})$ is the model’s text-image response, we first generate a set of three task-specific evaluation criteria $\mathcal{C}_{t}=\{c_{t,k}\}_{k=1}^{3}$ using GPT-5:
$c_{t,k}=\mathcal{G}_{\text{GPT-5}}(Q_{t},I_{t},T_{t}).$ (1)
@@
where $\mathcal{C}_{d}\subseteq\bigcup_{t=1}^{T}\mathcal{C}_{t}$ denotes the set of dynamically generated criteria assigned to dimension $d$. The assignment is based on syntactic and semantic cues in $Q_{t}$ and $T_{t}$. This results in a score triplet $\mathcal{S}=(\mathcal{S}_{\text{style}},\mathcal{S}_{\text{logic}},\mathcal{S}_{\text{entity}})\in[0,1]^{3}$, forming a multidimensional assessment of compositional performance. The protocol shows strong agreement with human judgment. We validate this by comparing model predictions with expert annotations on 3,000 held-out samples. The GPT-5-based evaluator achieves 87.6% agreement with human raters. This surpasses baseline methods by over 30 percentage points. Using static criteria with the same evaluator yields only 55.3% agreement.
@@ 5 Experiments
We conducted capability enhancement tests using IUT-Plug on existing interleaved VLMs input and output experimental benchmarks. We integrated the most advanced VLMs and text-to-image generation models (T2I models) to evaluate the information transfer capabilities across models. In this section, we aim to address the following key questions: (1) How do current interleaved VLMs perform in terms of understanding and generation across various benchmarks? (Such as MMIE (Xia et al., 2025))and OpenING (Zhou et al., 2025) (2) What are the distinctive features of IUT-Plug compared to existing methods, and in which aspects does it achieve the greatest improvements? For conclusions, we demonstrate IUT-Plug’s ability to improve key metrics.
@@ 5.1 Experimental Setup
Current interleaved VLMs can be categorized into three paradigms (Zhou et al., 2025): (1) Integrated pipelines, such as GPT-4 + DALL·E 3 and Gemini 1.5 + Flux (Achiam et al., 2023; OpenAI, 2023; Reid et al., 2024); (2) Two-stage generators; and (3) End-to-end generators. In this work, we focus on enhancing the integrated pipeline—a prominent approach within the interleaved VLM framework.
Text to image Models. Same as MMIE (Xia et al., 2025), we use three text-to-image models to integrate: (1) Openjourney (v4.1) (Community, 2024), a community-tuned Stable Diffusion variant excelling in artistic rendering; (2) SD-3 Medium (Esser et al., 2024), Stability AI’s flagship model optimized for photorealism and compositional accuracy; (3) Flux (1.0-dev) (Labs, 2024), an emerging architecture specializing in dynamic scene generation and spatial reasoning. These combinations create nine distinct VLM-T2I configurations for comprehensive benchmarking.
@@
Evaluation Models and Metrics. We fine-tuned Qwen2.5-VL-7B to construct an evaluation model using 3,000 expert-annotated samples. The annotation process involved domain experts scoring outputs across six dimensions. Specifically, we used GPT-5 and DALL·E to synthesize over 1,000 data samples, which were then scored by human experts from diverse domains across six dimensions (detailed scoring criteria. The full context, including reference scores, was used to fine-tune Qwen2.5-VL-7B.
Text to Image Set Evaluation Metrics.Since our method shows only marginal improvement on the macro-benchmark based on six metrics, we hypothesize that the enhancement might be specific to image-text coherence. Inspired by the approach of T2IS (Jia et al., 2025), we therefore propose the following benchmark.
@@
Finally, we use GPT-5 to generate dynamic evaluation criteria for assessing performance on these dimensions. We then employ Qwen3-235B to evaluate each image text QA pair against these criteria. For each criterion, the model outputs a “yes” or “no” response based on the provided instructions. We compute the logit probabilities of these responses and normalize them into a score between zero and one. This normalized value serves as the consistency score for that criterion. The final score for each dimension, entity style, or logic is the average of all corresponding criterion scores within that dimension. This process is illustrated in Figure 2.
@@ Appendix A Appendix
A.1 Key Components and Evaluation Criteria
@@
A.2 Overview of Baseline Models
@@
  • MiniGPT-5 (Zheng et al., 2023): Combines MiniGPT-4 and Stable Diffusion for multimodal I/O, using ”generative tokens” to bridge text and vision.
  • EMU-2 (Sun et al., 2024): A 37B parameter generative multimodal model. We use a pipeline of its chat (Emu2-Chat) and generation (Emu2-Gen) variants.
  • GPT-4o (Achiam et al., 2023): An advanced multimodal model from OpenAI capable of processing both text and visual inputs.
  • Gemini-1.5 (Reid et al., 2024): A large language model from Google AI trained on a massive text and code dataset, with capabilities for image analysis.
  • LLaVA-34b (Liu et al., 2023): An end-to-end model connecting a vision encoder with the Hermes-Yi-34B LLM for visual and language understanding. It does not support multiple image inputs.
  • Qwen2-VL-72b (Wang et al., 2024): The multimodal version of Alibaba Cloud’s Qwen large model series, designed for text, image, and audio processing.
@@
  • Stable Diffusion 3 Medium (Esser et al., 2024): A text-to-image model from Stability AI that generates high-quality images with fine detail.
  • Stable Diffusion XL turbo (Esser et al., 2024): An optimized version of SDXL for accelerated, high-quality image generation.
@@
A.3 Prompts for Image Generation Models
@@ A.4.1 IUT Performance Examples
A.4.2 Ablation Study Examples
@@ A.4.3 IUT Extraction Examples
A.5 Six-point grading system Criteria

### v2

@@ 1 Introduction
Despite remarkable advancements, modern Vision Language Models (VLMs) suffer from a fundamental limitation known as multimodal context drift. Multimodal context drift refers to a progressive loss of cross-modal consistency during extended interleaved image-text interactions. Although VLMs excel at interpreting static image-text pairs (Rombach et al., 2022; Ramesh et al., 2022), they often fail to maintain coherence across multi-turn sequences. Specifically, (1) Logic drift occurs when the joint semantic content derived from the input image and accompanying text is not faithfully reflected in either the textual answer or the subsequently generated image, leading to contradictions between perception, instruction, and output. (2) Entity identity drift manifests when concrete referents—such as named characters or objects present in the input image or text—are misidentified, lose their attributes, or disappear entirely in later generations. (3) Style drift happens when distinctive visual characteristics of the input image, including artistic medium, color palette, lighting, or composition, are not preserved in the output image despite being explicitly conditioned on that input (Malakouti and Kovashka, 2025; Said et al., 2025; Park et al., 2021). These three forms of drift reveal a systemic failure in current VLMs to maintain a unified, persistent representation of the multimodal scene across turns (Bougzime et al., 2025; Marcus, 2020).
@@
To bridge this gap, we propose IUT-Plug, a lightweight and model-agnostic plug-in module. IUT-Plug is built around Image Understanding Trees (IUTs), which are hierarchical symbolic structures. These structures explicitly represent the entities in a visual scene, their attributes, and the relationships among them. The module integrates seamlessly with existing large vision-language models without requiring costly end-to-end retraining (Li et al., 2024; Chen et al., 2024). It operates within the neuro-symbolic paradigm (d’Avila Garcez and Lamb, 2020; Sarker et al., 2021). By design, IUT-Plug separates the perceptual strength of neural networks from the logical precision of symbolic reasoning. It uses the IUT as a factual scaffold to guide both text and image generation (Ferreira et al., 2023; Pan et al., 2024). This combination allows the system to maintain consistency while preserving the flexibility of modern generative models.
@@
IUT-Plug bridges perception and generation. In the perception phase, the input image is parsed into an Image Understanding Tree (IUT), which represents entities, their attributes, and their relationships in a structured symbolic format. As new interleaved image-text instructions arrive, this state is incrementally updated to reflect evolving semantics. The refined IUT then actively guides both the language response and the prompt for the downstream text-to-image generator. By anchoring the entire generation process in this symbolic scaffold, IUT-Plug ensures that logical coherence, entity identity, and visual style are faithfully propagated from the input to all subsequent outputs. This design directly addresses the three core forms of multimodal context drift that undermine current interleaved VLM pipelines.
@@ 2.1 Interleaved Vision-Language Models
The ability to process and generate sequences of interleaved images and text marks a major advance over single turn multimodal systems. Pioneering architectures like Flamingo (Alayrac et al., 2022) introduced gated cross-attention to ingest multimodal streams. More recent systems have further enhanced generative fluency by integrating vision-language models (VLMs) with text-to-image (T2I) generators, such as MiniGPT (Zheng et al., 2023), Emu2 (Sun et al., 2024), and MM-Interleaved (Tian et al., 2024). They often using visual tokens or feature synchronizers to bridge modalities (Wu et al., 2024). However, no explicit mechanism exists to propagate the original visual context—especially its logic, entity identity, and style. As a result, the system suffers from multimodal context drift. The joint semantics of the input are progressively lost or distorted across turns.
@@ 2.2 Image Generation from Structured Representations
Structured symbolic representations have long been used to guide image synthesis. Much of this work centers on scene graphs (Chang et al., 2021). Models such as SG2IM (Johnson et al., 2018) and SGDiff (Yang et al., 2022) show that a static graph—encoding objects, their attributes, and pairwise relationships—can be effectively translated into a coherent image. The inverse task is addressed by Scene Graph Generation (SGG) models (Krishna et al., 2017), which parse a single image into such a graph. However, these approaches treat the structured representation as a one-time input or output. They do not support updates across interactions. As a result, they are inherently limited to single-turn generation. Multi-turn consistency remains out of reach in this paradigm.
@@ 2.3 Neuro-Symbolic Generative Models
Neuro-symbolic AI aims to combine the perceptual strength of neural networks with the reasoning power of symbolic systems (Marcus, 2020). In generative modeling, several methods have explored symbolic guidance. Neuro-Symbolic Diffusion (NSD) (Christopher et al., 2025) injects logical constraints directly into the denoising steps of diffusion models. Control-GPT (Zhang et al., 2023) uses programmatic sketches to control layout and composition. These approaches often rely on hard constraints or procedural specifications. While effective for constrained tasks, they can limit the generative flexibility of the underlying model. Moreover, they typically assume a fixed symbolic specification at the start of generation. They lack mechanisms to maintain or evolve a world state over extended, open-ended interactions.
@@ 3 Evaluation Framework
Figure 2: Overview of our evaluation metric. The evaluation model is fine-tuned on 3,000 sample data annotated by experts. For each question-answer pair, we use GPT-4o to generate three dynamic evaluation criteria, and then employ the evaluation model to output “yes” or “no” for each criterion.
@@
Seen as Figure 3. For each input-output tuple $(Q_{t},I_{t},A_{t})$, where $Q_{t}\in\mathcal{Q}$ is a natural language instruction, $I_{t}\in\mathbb{R}^{H\times W\times 3}$ is the reference image, and $A_{t}=(T_{t},I_{t}^{\prime})$ is the model’s text-image response, we first generate a set of three task-specific evaluation criteria $\mathcal{C}_{t}=\{c_{t,k}\}_{k=1}^{3}$ using GPT-4o:
$c_{t,k}=\mathcal{G}_{\text{GPT-4o}}(Q_{t},I_{t},T_{t}).$ (1)
@@
where $\mathcal{C}_{d}\subseteq\bigcup_{t=1}^{T}\mathcal{C}_{t}$ denotes the set of dynamically generated criteria assigned to dimension $d$. The assignment is based on syntactic and semantic cues in $Q_{t}$ and $T_{t}$. This results in a score triplet $\mathcal{S}=(\mathcal{S}_{\text{style}},\mathcal{S}_{\text{logic}},\mathcal{S}_{\text{entity}})\in[0,1]^{3}$, forming a multidimensional assessment of compositional performance. The protocol shows strong agreement with human judgment. We validate this by comparing model predictions with expert annotations on 3,000 held-out samples. The GPT-4o-based evaluator achieves 87.6% agreement with human raters. This surpasses baseline methods by over 30 percentage points. Using static criteria with the same evaluator yields only 55.3% agreement.
@@ 5 Experiments
We conducted capability enhancement tests using IUT-Plug on existing interleaved VLMs input and output experimental benchmarks. We integrated the most advanced VLMs and text-to-image generation models (T2I models) to evaluate the information transfer capabilities across models. In this section, we aim to address the following key questions: (1) How do current interleaved VLMs perform in terms of understanding and generation across various benchmarks? (Such as MMIE (Xia et al., 2024))and OpenING (Zhou et al., 2025) (2) What are the distinctive features of IUT-Plug compared to existing methods, and in which aspects does it achieve the greatest improvements? For conclusions, we demonstrate IUT-Plug’s ability to improve key metrics.
@@ 5.1 Experimental Setup
Current interleaved VLMs can be categorized into three paradigms (Zhou et al., 2025): (1) Integrated pipelines, such as GPT-4 + DALL·E 3 and Gemini 1.5 + Flux (Achiam et al., 2023; Betker et al., 2023; Team et al., 2024); (2) Two-stage generators; and (3) End-to-end generators. In this work, we focus on enhancing the integrated pipeline—a prominent approach within the interleaved VLM framework.
Text to image Models. Same as MMIE (Xia et al., 2024), we use three text-to-image models to integrate: (1) Openjourney (v4.1) , a community-tuned Stable Diffusion variant excelling in artistic rendering; (2) SD-3 Medium (Esser et al., 2024), Stability AI’s flagship model optimized for photorealism and compositional accuracy; (3) Flux (1.0-dev) (Labs et al., 2025), an emerging architecture specializing in dynamic scene generation and spatial reasoning. These combinations create nine distinct VLM-T2I configurations for comprehensive benchmarking.
@@
Evaluation Models and Metrics. We fine-tuned Qwen2.5-VL-7B to construct an evaluation model using 3,000 expert-annotated samples. The annotation process involved domain experts scoring outputs across six dimensions. Specifically, we used GPT-4o and DALL·E to synthesize over 1,000 data samples, which were then scored by human experts from diverse domains across six dimensions (detailed scoring criteria. The full context, including reference scores, was used to fine-tune Qwen2.5-VL-7B.
Text to Image Set Evaluation Metrics.Since our method shows only marginal improvement on the macro-benchmark based on six metrics, we hypothesize that the enhancement might be specific to image-text coherence. Inspired by the approach of T2IS (Anonymous, 2025), we therefore propose the following benchmark.
@@
Finally, we use GPT-4o to generate dynamic evaluation criteria for assessing performance on these dimensions. We then employ Qwen3-235B to evaluate each image text QA pair against these criteria. For each criterion, the model outputs a “yes” or “no” response based on the provided instructions. We compute the logit probabilities of these responses and normalize them into a score between zero and one. This normalized value serves as the consistency score for that criterion. The final score for each dimension, entity style, or logic is the average of all corresponding criterion scores within that dimension. This process is illustrated in Figure 2.
@@ Appendix A Appendix
A.1 Structured Text Prompting (STP) Baseline
The Structured Text Prompting (STP) baseline is introduced to rigorously validate the necessity of IUT-Plug’s persistent, incrementally updated symbolic state over merely using structured instructions within a single VLM turn. STP simulates an enhanced, state-of-the-art VLM capable of following complex formatting rules, yet lacks an external, memory-augmented reasoning module.
In the STP method, the base Vision-Language Model (VLM) is given the full history $(I_{0},Q_{1},\dots,Q_{t})$ and is explicitly instructed to generate its textual response $T_{t}$ followed by a structured, self-contained summary of the current scene state $S_{t}$. However, this state $S_{t}$ is not programmatically parsed, maintained, or updated by an external module; instead, the VLM must regenerate the entire description from scratch in the next turn, relying solely on its internal attention mechanisms to maintain long-term context.
Prompt template for the VLM in the STP baseline. (using $T_{t-1}$ as the previous model response and $Q_{t}$ as the current user instruction)
Instruction: Based on the image and the dialogue history provided, first generate your natural language response $T_{t}$. Following your response, you must generate a full JSON description of the current visual scene state, $S_{t}$. This JSON must list all entities, their attributes, and their relationships. JSON Schema (Must Follow):
  "scene_summary": "a short textual summary of the scene.",
  "style": {
    "artistic_medium": "...",
    "lighting": "...",
  "objects": [
    {"name": "...", "attributes": "...", "relationships": "... "}
    // ... all other key entities
  "relationships": [
    "entity_A [relation] entity_B",
    // ... all key relations
History and Input: [Previous Turns and Image History] Current Instruction ($Q_{t}$): [User Input $Q_{t}$]
The T2I model for STP is then prompted using the VLM’s natural language response $T_{t}$ concatenated with the VLM-generated JSON summary $S_{t}$, similar to how the full IUT-Plug module operates. The experimental results in Section 4.2 confirm that while STP offers an improvement over the vanilla baseline (proving that structured information helps), it consistently falls short of the IUT-Plug, demonstrating the critical role of the IUT’s explicit, incremental state update mechanism in mitigating multimodal context drift.
A.2 Key Components and Evaluation Criteria
@@
A.3 Overview of Baseline Models
@@
  • MiniGPT [Zheng et al., 2023]: Combines MiniGPT-4 and Stable Diffusion for multimodal I/O, using ”generative tokens” to bridge text and vision.
  • EMU-2 [Sun et al., 2024]: A 37B parameter generative multimodal model. We use a pipeline of its chat (Emu2-Chat) and generation (Emu2-Gen) variants.
  • GPT-4o [Achiam et al., 2023]: An advanced multimodal model from OpenAI capable of processing both text and visual inputs.
  • Gemini-1.5 [Team et al., 2024]: A large language model from Google AI trained on a massive text and code dataset, with capabilities for image analysis.
  • LLaVA-34b [Liu et al., 2023]: An end-to-end model connecting a vision encoder with the Hermes-Yi-34B LLM for visual and language understanding. It does not support multiple image inputs.
  • Qwen2-VL-72b [Wang et al., 2024]: The multimodal version of Alibaba Cloud’s Qwen large model series, designed for text, image, and audio processing.
@@
  • Stable Diffusion 3 Medium [Esser et al., 2024]: A text-to-image model from Stability AI that generates high-quality images with fine detail.
  • Stable Diffusion XL turbo [Esser et al., 2024]: An optimized version of SDXL for accelerated, high-quality image generation.
@@
A.4 Prompts for Image Generation Models
@@ A.5.1 IUT Performance Examples
A.5.2 Ablation Study Examples
@@ A.5.3 IUT Extraction Examples
A.6 Six-point grading system Criteria